# Strassen Attention: Unlocking Compositional Abilities in Transformers Based on a New Lower Bound Method

### v2

@@ Abstract
We propose a novel method to evaluate the theoretical limits of Transformers, allowing us to prove the first lower bounds against one-layer softmax Transformers with infinite precision. We establish those bounds for three tasks that require advanced reasoning. The first task, Match3 (Sanford et al., 2023), requires looking at all triples of positions. The second and third tasks address compositionality-based reasoning: one is composition of functions (Peng et al., 2024) and the other is composition of binary relations. We formally prove the inability of one-layer softmax Transformers to solve any of these tasks. In an attempt to overcome these limitations, we introduce Strassen attention and prove that with this mechanism a one-layer Transformer can in principle solve all these tasks. We also show that it enjoys sub-cubic running-time complexity, making it more scalable than similar previously proposed mechanisms, such as higher-order attention (Sanford et al., 2023). To complement our theoretical findings, we experimentally studied Strassen attention and compared it against standard (Vaswani et al, 2017), higher-order attention (Sanford et al., 2023) and triangular attention (Bergen et al. 2021). Our results help to disentangle all these attention mechanisms, highlighting their strengths and limitations. In particular, Strassen attention outperforms standard attention significantly on all the tasks. Altogether, understanding the theoretical limitations can guide research towards scalable attention mechanisms that improve the reasoning abilities of Transformers.
@@ 1 Introduction
Limitations of Transformers
@@
Lower bounds
A natural first step to explain why models struggle with some tasks is to study their expressivity— whether a given architecture is in-principle capable of solving the task in question. That is, whether there exists a choice of parameters that results in a model computing the function underlying the task solutions. If the answer is positive, then the chances are that the difficulty lies in efficiently learning the appropriate parameters. If the answer is negative, then a formal proof of this fact can be directly related to the poor performance observed in practice. Mathematical results of the latter type are referred to as lower bounds.
The first lower bounds were obtained for Transformers using hardmax attention [10]. Instead of computing attention as a convex combination of all input tokens using softmax, one takes a single token where the attention is maximal. Using this simplification, Hahn showed that hardmax Transformers with $O(1)$ layers cannot compute formal languages such as PARITY, MAJORITY, or Dyck-1.
Lower bounds against softmax Transformers have recently been obtained by employing a different proof technique, one based on communication complexity [20]. The idea is to show that a Transformer solving a given task can be used to construct a corresponding communication protocol for a problem whose communication complexity is known. From this, one obtains lower bounds on the size of Transformers capable of solving the task for inputs of a given length. In Peng et al. [20] for example, the authors apply this technique to show that any one-layer softmax Transformer that can compute the composition of two functions must have $n^{\Omega(1)}$ embedding dimension, where $n$ is the size of the input. Crucially, for this conclusion to hold, one must assume that the Transformer works with a relatively low number of precision bits, namely sub-linear in $n$. The technique has subsequently been applied to show lower bounds for other tasks such as string equality [3] and Match3 [21] (see Section 3.2.2 for a definition).
@@
Equipped with a better understanding of these limitations, the next question is how they can guide research towards the construction of more expressive machines. A key observation is that the standard attention mechanism can only see interactions between pairs of tokens, whereas compositional tasks require to reason about the interaction of three or more tokens simultaneously [2, 21]. Consequently, suitably modified attention mechanisms have been proposed, which would track interactions between more than two tokens. For instance, triangular attention [2] (see Section 2 for a definition) outperforms standard attention in compositional tasks such as CLUTRR [23] or CoGS [13]. Similarly, it has been shown that higher-order tensor attention embedded in a constant size one-layer Transformer can theoretically solve Match3, a task that cannot be solved by the standard attention [21]. However, both these mechanisms suffer from a significant increase in running-time complexity, e.g., order 3 requires cubic-time, limiting its scalability and potentially affecting its practical relevance. In the case of triangular attention, the mechanism works with an adjacency matrix of a graph (in which case it is square-time in the size of the graph). It is possible to produce an adjacency matrix from the sequence of tokens, as already shown in Bergen et al. [2], but this increases the complexity to cubic. The higher-order tensor attention, on top of being cubic-time as well, has been only considered in theoretical work and has not been evaluated empirically yet.
@@
[…]
@@
    Main Theorem. Let $T$ be any one-layer standard attention Transformer that computes $f$. Denote by $d$, $H$ and $L$, respectively, its embedding dimension, number of attention heads and size of the output MLP of $T$. Then it holds that $\max\{d,H,L\}\geq\textsf{split-VC}(f)^{\Omega(1)}$.
    Instead of relying on communication complexity, our technique employs upper bounds on the VC-dimension of hypothesis classes of functions that can be computed with a limited number of elementary operations [9]. Remarkably, since communication complexity is not involved, our bounds do not need to assume any limit on the number of precision bits and apply even in the infinite precision regime.
[…]
@@
  • We develop a more scalable higher-order mechanism. In an attempt to address the scalability issues of previously introduced 3-order mechanisms, we present and study Strassen attention, a variation devised to be sensitive to interactions between triplets of tokens without sacrificing too much efficiency. It can be implemented in the form of $O(1)$ products of $n\times n$ matrices for inputs of size $n$. It thus enjoys $n^{2}$ space and sub-cubic running time due to fast matrix multiplication algorithms [^2] The first such algorithm was invented by Strassen [24], inspiring the name of our attention mechanism., making it more scalable than previously proposed mechanism with similar capacity. We formally prove that each of the four aforementioned tasks can be solved in-principle with a one-layer Strassen attention transformer of constant size.
@@
  • We demonstrate the empirical value of Strassen attention. To substantiate our theoretical findings, we experimentally tested our proposed mechanism on all these tasks. The results show that our attention systematically: (i) outperforms or matches the accuracy of other attention mechanisms, and (ii) achieves superior efficiency in terms of training time and computational resources when compared with other triple-wise interaction attention mechanisms. We compared Strassen’s performance against standard, triangular, and third-order attention, focusing on both accuracy and time required for the training process. By analyzing how quickly each method reached solutions, we obtain empirical validation that Strassen attention offers significant advantages in scalability.
@@
In Section 2 we recall the Transformer architecture and introduce our new Strassen attention mechanism. Section 3 presents our new lower bound method and its application to function composition, Match3, and binary relation composition tasks. In Section 4, we show that Strassen attention can be implemented in sub-cubic time, and formally prove that, in principle, it is capable of solving the three aforementioned task. Section 5 presents a novel task allowing to separate the capabilities of Strassen attention from those of standard and triangular attentions. Finally, section 6 contains our experimental findings. Due to space constraints, some proofs and experimental details are deferred to the Appendix.
@@ 2 Preliminaries
Throughout the paper, we denote $[n]=\{1,\ldots,n\}$ for $n\in\mathbb{N}$. For a set $\Sigma$, we will denote by $\Sigma^{n}$ the collection of sequences of elements of $\Sigma$ of length $n$, and by $\Sigma^{*}$ the collection of all finite sequences. We start by briefly recalling the basics of the Transformer architecture and formally defining the attention mechanisms studied in this paper.
The main block of the Transformer layer is the attention function, formally defined as a length-preserving function $a\colon(\mathbb{R}^{d})^{*}\to(\mathbb{R}^{d})^{*}$, where $d$ is the embedding dimension. In this paper, we consider 4 types of attention functions (Figure 1).
@@
where $W^{f},W^{g},W^{h},V_{1},V_{2}\in\mathbb{R}^{d\times d}$, and $\odot$ denotes the Hadamard product.
@@
A one-layer Transformer $T$ with $H$ heads and embedding dimension $d$ is given by
  • $H$ attention functions $Att_{1},\ldots,Att_{H}\colon(\mathbb{R}^{d})^{*}\to(\mathbb{R}^{d})^{*}$
  • a matrix $W_{O}\in\mathbb{R}^{d\times(dH)}$;
  • an “output MLP” $\mathcal{N}\colon\mathbb{R}^{d}\to\mathbb{R}$ with $P$ parameters, which is formally a neural network with ReLU activation.
We define the size of $T$ as $size(T)=\max\{H,d,P\}$. The output of $T$ on input $\bar{x}=(x_{1},\ldots,x_{n})\in(\mathbb{R}^{d})^{n}$ is the sequence $\bar{y}=(y_{1},\ldots,y_{n})\in(\mathbb{R})^{n}$ given by
@@
3 Lower bounds via Split-VC dimension
@@
Consider a function $f\colon\Sigma^{n}\to\{0,1\}$, where $\Sigma$ is a finite set of symbols. For a set of positions $A\subseteq\{1,\ldots,n\}$, we define a Boolean matrix $M_{f}^{A}$ as follows. Its rows will be indexed by all the words $w^{1}\in\Sigma^{A}$ and its columns by the words $w^{2}\in\Sigma^{B}$, where $B=\{1,\ldots,n\}\setminus A$. Thus, $M_{A}^{f}$ will be a $|\Sigma|^{|A|}\times|\Sigma|^{n-|A|}$ Boolean matrix. The value of $M_{A}^{f}$ at $(w^{1},w^{2})$ is then defined as
@@
where $w^{1}\oplus w^{2}\in\Sigma^{n}$ is obtained by merging $w^{1}$ and $w^{2}$ according to the positions indicated by $A$ and $B$.
@@
[…]
@@ 3.1 Main Theorem
On the other hand, the VC dimension of $H$ can be upper bounded by $m^{o(1)}$. Indeed, the number of parameters of $H$ is $dH+H=m^{o(1)}$. To compute the value of a given hypothesis on a given input, it is enough to do $m^{o(1)}$ basic arithmetic operations and comparisons with 00, because the size of $\mathcal{N}$ is $m^{o(1)}$. By Theorem 2.3 in [9], the VC dimension is polynomial in these quantities, which gives us $m^{o(1)}$ upper bound in our case. ∎
@@
[…]
@@ 3.2 Applications to three concrete tasks
We now apply Theorem 3.2 to three different tasks: function composition, Match3 and binary relation composition.
@@ 3.2.1 Function Composition
Introduced in Peng et al. [20], in this task we receive a description of two functions $g:[n]\to[n]$ and $h:[n]\to[n]$, and we are asked the value of $h(g(x))$ for a given $x\in[n]$. The task is presented to a Transformer in the form of a sequence of $12n+$ 1 tokens, divided in three parts. The first two parts encode the values of $g$ and $h$, and the third part encodes $x$ in a single token, where the output has to be computed. More specifically, the output has to be a real number $y_{2n+1}$ satisfying $|y_{2n+1}-h(g(x))|<0.5$. That is, with $h(g(x))$ being the closest integer to $y_{2n+1}$.
@@
$Ind_{n}(p_{1},q_{1},\ldots,q_{n})$ =$=\begin{cases}1&q_{p_{1}}=1,\\ 0&\text{otherwise,}\end{cases}$

### v3

# Strassen Attention, Split VC Dimension and Compositionality in Transformers

@@ Abstract
We propose the first method to show theoretical limitations for one-layer softmax transformers with arbitrarily many precision bits (even infinite). We establish those limitations for three tasks that require advanced reasoning. The first task, Match 3 (Sanford et al., 2023), requires looking at all possible token triplets in an input sequence. The second and third tasks address compositionality-based reasoning: function composition (Peng et al., 2024) and binary relations composition, respectively. We formally prove the inability of one-layer softmax Transformers to solve any of these tasks. To overcome these limitations, we introduce Strassen attention and prove that, equipped with this mechanism, a one-layer transformer can in principle solve all these tasks. Importantly, we show that it enjoys sub-cubic running-time complexity, making it more scalable than similar previously proposed mechanisms, such as higher-order attention (Sanford et al., 2023). To complement our theoretical findings, we experimentally studied Strassen attention and compared it against standard (Vaswani et al, 2017), higher-order attention (Sanford et al., 2023), and triangular attention (Bergen et al. 2021). Our results help to disentangle all these attention mechanisms, highlighting their strengths and limitations. In particular, Strassen attention outperforms standard attention significantly on all the tasks. Altogether, understanding the theoretical limitations can guide research towards scalable attention mechanisms that improve the reasoning abilities of Transformers.
@@ 1 Introduction
What tasks can Transformers solve?
@@
Theoretical limitations of Transformers
A natural first step to explain why models struggle with some tasks is to study their expressivity— whether a given architecture is in-principle capable of solving the task in question. That is, whether there exists a choice of parameters that results in a model computing the function underlying the task solutions. If the answer is positive, then the chances are that the difficulty lies in efficiently learning the appropriate parameters. If the answer is negative, then a formal proof of this fact can be directly related to the poor performance observed in practice.
The first theoretical limitations of this sort were obtained for Transformers using hardmax attention [10]. Instead of computing attention as a convex combination of all input tokens using softmax, one takes a single token where the attention is maximal. Using this simplification, Hahn showed that hardmax Transformers with $O(1)$ layers cannot compute formal languages such as PARITY, MAJORITY, or Dyck-1.
Theoretical limitations against softmax Transformers have recently been obtained by employing a different proof technique, one based on communication complexity [19]. The idea is to show that a Transformer solving a given task can be used to construct a corresponding communication protocol for a problem whose communication complexity is known. From this, one obtains lower bounds on the size of Transformers capable of solving the task for inputs of a given length. In [19], for example, the authors apply this technique to show that any one-layer softmax Transformer that can compute the composition of two functions must have $n^{\Omega(1)}$ embedding dimension, where $n$ is the size of the input. Crucially, for this conclusion to hold, one must assume that the Transformer works with a relatively low number of precision bits, namely sub-linear in $n$. The technique has subsequently been applied to show lower bounds for other tasks such as string equality [3] and Match3 [20] (see Section 3.2.2 for a definition).
@@
Equipped with a better theoretical understanding of why Transformers struggle with some tasks, the next question is how they can guide research towards the construction of more expressive machines. A key observation is that the standard attention mechanism can only see interactions between pairs of tokens, whereas compositional tasks require to reason about the interaction of three or more tokens simultaneously [2, 20]. Consequently, suitably modified attention mechanisms have been proposed, which would track interactions between more than two tokens. For instance, triangular attention [2] (see Section 2 for a definition) outperforms standard attention in compositional tasks such as CLUTRR [22] or CoGS [13]. Similarly, it has been shown that higher-order tensor attention embedded in a constant size one-layer Transformer can theoretically solve Match3, a task that cannot be solved by the standard attention [20]. However, both these mechanisms suffer from a significant increase in running-time complexity, e.g., order 3 requires cubic-time, limiting its scalability and potentially affecting its practical relevance. In the case of triangular attention, the mechanism works with an adjacency matrix of a graph (in which case it is square-time in the size of the graph). It is possible to produce an adjacency matrix from the sequence of tokens, as already shown in Bergen et al. [2], but this increases the complexity to cubic. The higher-order tensor attention, on top of being cubic-time as well, has been only considered in theoretical work and has not been evaluated empirically yet.
@@
  • We present a new technique for proving theoretical limitations. Since previous results establishing limitations for softmax transformers are based on communication complexity, they only apply to transformers that store numbers with relatively few precision bits. This raises the question – would the use of long arithmetic help to circumvent these limitations? In this paper, we answer this question negatively. In order to establish that, we introduce a new technique for proving limitations, based on the novel notion of Splitting VC dimension.
    The splitting VC dimension of a Boolean function $f$, denoted by $\textsf{split-VC}(f)$, is a positive integer number intended to capture the complexity of $f$ in a certain combinatorial sense. Our main technical contribution is the following:
@@
    Main Theorem. Let $T$ be any one-layer standard attention Transformer that can do arithmetic operations over real numbers with infinite precision. Denote by $d$, $H$ and $L$, respectively, its embedding dimension, number of attention heads and size of the output MLP of $T$. Suppose $T$ can compute a function $f$. Then it holds that $\max\{d,H,L\}\geq\textsf{split-VC}(f)^{\Omega(1)}$.
    To put it differently, even an idealized transformer $T$ that can manipulate numbers with infinitely many bits cannot compute a function $f$ with large $\textsf{split-VC}(f)$ unless $T$ has a large embedding dimension, a large number of attention heads, or a large output MLP. Obviously, the same limitations apply to real-life Transformers that can only manipulate numbers using some finite (however large) number of bits.
  • We obtain new theoretical limitations for tasks requiring complex reasoning. We apply our new method to the task of function composition (Peng et al. [19]) and the Match3 (Sanford et al. [20]) task, and show that previously known limitations for these tasks are also applicable to transformers that work with arbitrarily large (and even infinite) precision.
    We also do this for a new task, that we name binary relation composition task. We introduce it because the function composition task can be solved by a 2-layer transformer, while for the binary relation composition task there is no apparent solution with any constant number of layers. Finally, we introduce another task that we call Quotient Binary Relation Composition (see Section 5). Our method allows to establish limitations for this task even for mixed transformers that can use both the standard and the triangular attention.
@@
  • We develop a more scalable higher-order mechanism. Previous works (e.g. [20]) have shown that higher-order methods can in principle overcome these limitations, but suffer from severe scalability issues. In an attempt to address these difficulties, we present and study Strassen attention, a variation devised to be sensitive to interactions between triplets of tokens without sacrificing too much efficiency. In contrast to previous similar attention mechanisms that required running time $n^{3}$ for inputs of length $n$, our mechanism enjoys $n^{2}$ space and sub-cubic running time, making it more scalable. At the same time, we show that 1-layer constant-size Strassen attention transformers are theoretically capable of solving all 4 aforementioned tasks.
@@
  • We empirically demonstrate the convergence of Strassen attention. Is Strassen attention capable of converging to these theoretical solutions during learning? We empirically demonstrate that the answer is positive for all 4 tasks. For comparison, we have additionally evaluated all other attention mechanisms on these tasks. Our empirical findings agree with theory and show that Strassen attention: (i) outperforms the accuracy of Standard attention, (ii) achieves superior efficiency in terms of training time and computational resources when compared with other triple-wise interaction attention mechanisms.
@@
In Section 2 we recall the Transformer architecture and introduce our new Strassen attention mechanism. Section 3 presents our new lower bound method and its application to function composition, Match3, and binary relation composition tasks. In Section 4, we show that Strassen attention can be implemented in sub-cubic time, and formally prove that, in principle, it is capable of solving the three aforementioned task. Section 5 presents a novel task allowing to separate the capabilities of Strassen attention from those of standard and triangular attentions. Finally, section 6 contains our experimental findings.
@@ 2 Preliminaries
Throughout the paper, we denote $n]=\{1,\ldots,n\}$ for $n\in\mathbb{N}$. For a set $\Sigma$, we will denote by $\Sigma^{n}$ the collection of sequences of elements of $\Sigma$ of length $n$, and by $\Sigma^{*}$ the collection of all finite sequences. We start by briefly recalling the basics of the Transformer architecture and formally defining the attention mechanisms studied in this paper.
The main block of the Transformer layer is the attention function, formally defined as a length-preserving function $a\colon(\mathbb{R}^{d})^{*}\to(\mathbb{R}^{d})^{*}$, where $d$ is the embedding dimension. In this paper, we consider 4 types of attention functions.
@@
where $W^{f},W^{g},W^{h},V_{1},V_{2}\in\mathbb{R}^{d\times d}$, and $\odot$ denotes the Hadamard product. See Figure 1 for the illustration of these attention mechanisms.
@@
A one-layer Transformer $T$ with $H$ heads and embedding dimension $d$ is given by $H$ attention functions $Att_{1},\ldots,Att_{H}\colon(\mathbb{R}^{d})^{*}\to(\mathbb{R}^{d})^{*}$, a matrix $W_{O}\in\mathbb{R}^{d\times(dH)}$, and an “output MLP” $\mathcal{N}\colon\mathbb{R}^{d}\to\mathbb{R}$ with $P$ parameters, which is formally a neural network with ReLU activation. We define the size of $T$ as $size(T)=\max\{H,d,P\}$. The output of $T$ on input $\bar{x}=(x_{1},\ldots,x_{n})\in(\mathbb{R}^{d})^{n}$ is the sequence $\bar{y}=(y_{1},\ldots,y_{n})\in(\mathbb{R})^{n}$ given by
@@
3 Theoretical Limitations of Transformers via Split-VC dimension
@@
We now explain how to adapt VC dimension to use it as a complexity measure of a single function (instead of a class of functions). Take a function $f\colon\Sigma^{n}\to\{0,1\}$ and imagine we split the $n$ arguments of $f$ into two parts. One part will continue to correspond to the inputs, but the other is to be regarded as a set of parameters, so that now we can see $f$ as a class of functions that has a well-defined VC dimension. The complexity of $f$ will be defined as the maximal VC dimension that we can obtain in this way, considering all possible splittings into parameters and inputs.
Let us formalize this idea. For a set of positions $A\subseteq\{1,\ldots,n\}$, we define a Boolean matrix $M_{f}^{A}$ as follows. Its rows (interpreted as inputs) will be indexed by all the words $w^{1}\in\Sigma^{A}$ and its columns (interpreted as parameters) by the words $w^{2}\in\Sigma^{B}$, where $B=\{1,\ldots,n\}\setminus A$. Thus, $M_{A}^{f}$ will be a $|\Sigma|^{|A|}\times|\Sigma|^{n-|A|}$ Boolean matrix. The value of $M_{A}^{f}$ at $(w^{1},w^{2})$ is then defined as
@@
where $w^{1}\oplus w^{2}\in\Sigma^{n}$ is obtained by merging $w^{1}$ and $w^{2}$ according to the positions indicated by $A$ and $B$, i.e.
$(w^{1}\oplus w^{2})_{i}=\begin{cases}w^{1}_{i}&i\in A,\\ w^{2}_{i}&i\in B,\end{cases}\qquad i=1,\ldots,n.$
@@
We define the splitting VC dimension of $f\colon\Sigma^{n}\to\{0,1\}$, denoted by $\textsf{split-VC}(f)$, as the maximum over $A\subseteq\{1,\ldots,n\}$ of the VC dimension of the set of columns of $M_{f}^{A}$, understood as Boolean functions on the set of rows.
Example.
Consider a Boolean function $f\colon\{0,1\}^{4}\to\{0,1\}$, defined by $f(x_{1},x_{2},x_{3},x_{4})=(x_{1}\land x_{2})\oplus(x_{3}\land x_{4})$. We claim that $\textsf{split-VC}(f)=2$. To establish $\textsf{split-VC}(f)\geq 2$, we consider $A=\{1,3\}$. The matrix $M_{f}^{A}$ is constructed as follows: its rows are indexed by Boolean vectors $x_{1}x_{3}\in\{0,1\}^{2}$, its columns by Boolean vectors $x_{2}x_{4}\in\{0,1\}^{2}$, and the intersection of the row $x_{1}x_{3}$ and column $x_{2}x_{4}$ has $f(x_{1},x_{2},x_{3},x_{4})$ in it. Thus, this matrix looks as follows:
$x_{1}x_{3}$ $x_{2}x_{4}$ 00 01 10 11
00 0 0 0 0
01 0 1 0 1
10 0 0 1 1
11 0 1 1 0
Columns of this matrix realize all 4 Boolean functions on the second and the third row, implying that the VC dimension of the set of columns of this matrix is at least 2. We now observe that there is no $A\subseteq\{1,2,3,4\}$ such that the set of columns of $M_{f}^{A}$ has VC dimension at least 3. Indeed, this requires $A$ to be of size at least 2, because otherwise there are less than 3 rows. But if $|A|\geq 2$, there are at most $4$ columns, making it impossible to realize $2^{3}=8$ different Boolean functions.
@@ 3.1 Main Theorem
On the other hand, the VC dimension of $H$ can be upper bounded by $m^{o(1)}$. Indeed, the number of parameters of $H$ is $dH+H=m^{o(1)}$. To compute the value of a given hypothesis on a given input, it is enough to do $m^{o(1)}$ basic arithmetic operations and comparisons with 0, because the size of $\mathcal{N}$ is $m^{o(1)}$. By Theorem 2.3 in [9], the VC dimension is polynomial in these quantities, which gives us $m^{o(1)}$ upper bound in our case. ∎
@@
Note that from Theorem 3.2 it follows that for any Transformer $T$ satisfying $size(T)=n^{o(1)}$, it is impossible to compute in an auxiliary token any function $f\colon\Sigma^{n}\to\{0,1\}$ with $\textsf{split-VC}(f)=n^{\Omega(1)}$.
@@ 3.2 Applications to three concrete tasks
We now apply Theorem 3.2 to three different tasks.
@@ 3.2.1 Function Composition
Introduced in Peng et al. [19], in this task we receive a description of two functions $g:[n]\to[n]$ and $h:[n]\to[n]$, and we are asked the value of $h(g(x))$ for a given $x\in[n]$. The task is presented to a Transformer in the form of a sequence of $2n+1$ tokens, divided in three parts. The first two parts encode the values of $g$ and $h$, and the third part encodes $x$ in a single token, where the output has to be computed. More specifically, the output has to be a real number $y_{2n+1}$ satisfying $|y_{2n+1}-h(g(x))|<0.5$. That is, with $h(g(x))$ being the closest integer to $y_{2n+1}$.
@@
$Ind_{n}(p_{1},q_{1},\ldots,q_{n})$ $=\begin{cases}1&q_{p_{1}}=1,\\ 0&\text{otherwise,}\end{cases}$